From an inventory of 48 new cars being shipped to local dealerships, corporate
reports indicate that 12 have defective radios installed. What is the probability that out of 8 new cars it just received, a dealership would discover, after testing, that exactly half of the cars have defective radios?

Answer by stanbon(75887)

From an inventory of 48 new cars being shipped to local dealerships, corporate
reports indicate that 12 have defective radios installed.
What is the probability that out of 8 new cars it just received, a dealership would discover, after testing, that exactly half of the cars have defective radios?
Binomial Problem with n = 8 and p = 12/48 = 1/4
------
P(x = 4) = 8C4(0.25)^4*(0.75)^4 = 0.08652
---------------
Cheers,
Stan H.